# Estimation and goodness-of-fit testing for positive random variables with explicit Laplace transform

### v1

@@ 7 Discussion and concluding remarks
In this paper, we proposed a new general inferential approach based on the Laplace transform of a random variable, that is based on exponential random censoring. This approach is useful to derive tractable parameter estimators that are used to propose suitable GOF tests for some non-standard families of random variables. The proposed estimators and test statistics are proven to be asymptotically normal leading to computationally convenient inferential procedures. We show how exponential censoring can be used to build parameter estimators and GOF tests for the positive stable and for the Tweedie laws for which specific tests are not present in the literature. The simulation study shows that the finite-sample performance of the proposed procedure is rather satisfactory even in the three-parameter case for a moderately large sample size.

### v2

@@ 7 Discussion and concluding remarks
We propose a new general inferential approach based on the Laplace transform of a random variable, that relies on exponential random censoring. This approach is useful to derive tractable parameter estimators that are used to propose suitable GOF tests for some non-standard families of random variables. The proposed estimators and test statistics are proven to be asymptotically normal, leading to computationally convenient inferential procedures. We show how exponential censoring can be used to build parameter estimators and GOF tests for the positive stable and for the Tweedie laws for which specific tests are not present in the literature. The simulation study shows that the finite-sample performance of the proposed procedures is rather satisfactory even in the three-parameter case for a moderately large sample size.